# Fidel-TS: A High-Fidelity Benchmark for Multimodal Time Series Forecasting

### v2

@@ 4.2 Experimental Setup
(1) Unimodal Forecasting Models, including MLP-based Dlinear (Zeng et al., 2023), FITS (Xu et al., 2024b), transformer-based PatchTST (Nie et al., 2023), iTransformer (Liu et al., 2024b), LLM-based GPT4TS (Zhou et al., 2023), and advanced foundation models Chronos (Ansari et al., 2024), TimeMoE (Shi et al., 2025) and Sundial (Liu et al., 2025), which pre-trained on vast numerical time series corpora.
(2) Multimodal Forecasting Models, including LLM-based GPT4MTS (Jia et al., 2024) and transformer-based FIATS (Xu et al., 2024a). They have different input requirements due to their distinct modeling approaches.
(3) LLMs, including general-purpose LLMs Qwen-2.5-14B-Instruct (Yang et al., 2024), Qwen-2.5-14B-Instruct-1M (Yang et al., 2025b), Qwen-3-14B (Yang et al., 2025a), Deepseek-R1 (DeepSeek-AI et al., 2025), and expert LLM Chattime (Wang et al., 2025) which fine-tuned based on Llama-2-7b (Touvron et al., 2023). To accommodate LLMs for forecasting, we designed prompts for both unimodal and multimodal forecasting tasks.
@@
Table 2: Performance of unimodal forecasting models on the full test set of each dataset. All metrics are Mean Squared Error (MSE). The best results are highlighted in red and the second-best in blue. Domain-specific models are trained on the training set of each dataset. Foundation models are evaluated in zero-shot forecasting without fine-tuning on each training set.
Dataset
Pred.
Len.
Domain-Specific Foundation Models
PatchTST iTrans. Dlinear FITS
GPT4TS
Chronos
Time-MoE
Sundial
@@
Realistic Settings
To ensure our evaluation mirrors the practical scenario of forecasting, we extend our high-fidelity principle from the benchmark to the experimental setup itself. To simulate diverse real-world time scales, our design aligns forecasting horizons with intuitive and operational periods. For instance, for datasets with an hourly frequency, we employ historical horizon of 360 (indicates half a month) and prediction lengths of {24, 168, 336, 720} which correspond to {1 day, 1 week, 2 weeks, and 1 month}. Our framework not only supports classic ratio-based splitting (e.g., 7:1:2) (Zhou et al., 2021; Wu et al., 2021), but also enables splitting by timestamps (e.g. Jan. 1st, 2021, Jan. 1st, 2022). This approach makes the evaluation setup more transparent and controllable, providing further safeguard against potential leakage issue. See Appendix D for more details of settings.
4.3 Benchmarking Unimodal Forecasting
@@
The evaluation results of all unimodal forecasting models on Fidel-TS are presented in Table 2. The results show that: (1) Although most of the domain-specific models have previously claimed state-of-the-art status on classic unimodal benchmarks, the performance of them is highly dataset-dependent on our benchmark, with no single model consistently emerging as a front-runner. (2) while large-scale foundation models claim to "fit all time series" and can outperform domain-specific models in zero-shot forecasting, their performance is merely comparable to specialized models on short forecast horizons and degrades significantly over longer ones. These suggest that the improvements observed on classic unimodal benchmarks are likely biased and do not generalize, questioning their validity in assessing the true capability of the model. See more results and analysis in Appendix E.1.
Furthermore, classic benchmarks often employ unreasonable transfer strategies (Jin et al., 2024), forcing models to transfer between datasets with entirely different domain (e.g., ETT $\rightarrow$ Weather), or between subsets of the same domain with mismatched sampling frequencies (e.g., ETTh1 $\rightarrow$ ETTm1). In contrast, our distinction between subjects and channels enables a more realistic transfer setup. We train models exclusively on the -obs subsets of NYTS and BEAR and evaluate them directly on the held-out -hid subsets, simulating a more practical scenario.
@@
Table 3: Generalization test of domain-specific unimodal forecasting models. All metrics are MSE. The best results are highlighted in red and the second-best in blue.
Dataset
Pred.
Len.
PatchTST iTrans. Dlinear FITS GPT4TS
NYTS-h-obs
$\rightarrow$ NYTS-h-hid
24 0.611 0.606 0.650 0.655 0.603
168 1.123 1.142 1.187 1.228 1.162
336 1.310 1.329 1.374 1.440 1.325
720 1.424 1.445 1.501 1.602 1.513
@@
$\rightarrow$ BEAR-hid
12 0.100 0.059 0.105 0.139 0.094
144 0.424 0.382 0.456 0.507 0.426
288 0.588 0.549 0.595 0.659 0.579
576 0.762 0.717 0.801 0.906 0.876
As presented in Table 3, the performance of the models is inconsistent, mirroring the results in Table 2. In our generalization test which is designed to be more meaningful in real-world scenarios, no single model demonstrates superior transfer capabilities. This suggests that previous transfer strategies may also introduce bias, producing an illusion of strong transferability in unrealistic tasks, yet this effect does not hold in practical, real-world scenarios.
Table 4: Performance of LLMs (for unimodal forecasting) on -mini test set of each dataset. All metrics are MSE; for LLMs, we additionally report the pass rate within 3 allowed attempts. The best results are highlighted in red and the second-best in blue.
Dataset
Pred.
Len.
LLMs (Unimodal forecasting) Domain-Specific Foundation Models
Qwen2.5
-14B
Instruct
Qwen2.5
-14B
Instruct
-1M
Qwen3
-14B
Chattime
Deepseek
-R1
PatchTST iTrans. Dlinear FITS GPT4TS Chronos
Time
-MoE
Sundial
128K 1M 128K 4K 128K
CCP-mini 24
0.311
@96.3%
0.289
@96.9%
0.235
@98.8%
0.263
@100.0%
0.182
@99.9%
0.140 0.127 0.138 0.141 0.129 0.144 0.146 0.131
168
0.865
@75.0%
0.374
@73.6%
0.298
@90.6%
0.457
@100.0%
0.267
@97.5%
0.192 0.169 0.197 0.197 0.179 0.203 0.219 0.192
GREG-mini 24
0.229
@96.9%
0.213
@97.6%
0.153
@96.7%
0.223
@100.0%
0.113
@99.8%
0.073 0.070 0.089 0.086 0.090 0.093 0.089 0.085
168
0.497
@74.4%
0.279
@73.2%
0.263
@88.0%
0.486
@100.0%
0.213
@93.2%
0.135 0.135 0.150 0.137 0.143 0.170 0.162 0.154
JAP-mini 24
@0.0%
1.825
@29.1%
@0.0%
0.483
@100.0%
@0.0%
0.286 0.285 0.334 0.334 0.278 0.277 0.278 0.286
168
@0.0%
@0.0%
@0.0%
0.580
@100.0%
@0.0%
0.418 0.411 0.419 0.429 0.421 0.443 0.449 0.448
NYTS-mini 24
1.884
@93.2%
1.893
@93.7%
2.236
@97.7%
0.998
@100.0%
1.071
@99.9%
0.826 0.764 0.864 0.871 0.774 0.880 0.923 0.777
168
3.599
@72.9%
2.378
@69.4%
2.546
@90.0%
2.312
@100.0%
2.154
@95.3%
1.538 1.599 1.560 1.615 1.749 1.876 1.675 1.797
CAISO-mini 24
@0.0%
0.501
@45.8%
@0.0%
0.216
@100.0%
@0.0%
0.136 0.140 0.143 0.145 0.137 0.146 0.132 0.130
168
@0.0%
@0.0%
@0.0%
0.453
@100.0%
@0.0%
0.276 0.297 0.324 0.285 0.278 0.297 0.298 0.282
BEAR-mini 12
0.944
@96.7%
0.522
@95.2%
4.279
@99.8%
0.145
@100.0%
0.299
@99.9%
0.103 0.080 0.111 0.144 0.091 0.126 0.118 0.116
144
2.205
@11.6%
7.425
@58.1%
1.486
@62.8%
1.077
@100.0%
2.249
@88.4%
0.484 0.353 0.505 0.593 0.432 0.836 0.831 1.027
Moreover, the evaluation results of LLMs for unimodal forecasting (combined with the test results of unimodal forecasting models on -mini datasets) are summarized in Table 4. The results show that despite prior claims of performance enhancement through prompting (Gruver et al., 2023), LLMs consistently underperform specialized unimodal models in all tested scenarios on our benchmark. This suggests that when provided solely with historical time series as input, LLMs struggle to effectively capture temporal patterns and extrapolate numerical values for forecasting. The improvements observed in previous studies are likely artifacts of flawed evaluation, stemming from either pre-training contamination on outdated benchmarks, which allows LLMs to memorize answers, or overly short prediction horizons that render the task trivially simple.
4.4 Benchmarking Multimodal Forecasting models and LLMs
Table 5: Performance of multimodal forecasting models and LLMs (for multimodal forecasting) on -mini test set of each dataset. All metrics are MSE; for LLMs, we additionally report the pass rate within 3 attempts. We first use control events for the BEAR dataset. The best results are highlighted in red and the second-best in blue.
Dataset Pred. LLMs (Multimodal forecasting) Domain-Specific
Qwen2.5-14B
-Instruct
Qwen2.5-14B
-Instruct-1M
Qwen3-14B
Chattime
Deepseek-R1
GPT4MTS FIATS
128K 1M 128K 4K 128K
CCP-mini 24
0.315
@96.8%
0.303
@97.1%
0.222
@95.6%
0.313
@100.0%
0.137 @99.8% 0.146 0.093
168
0.570
@71.4%
0.555
@81.5%
0.336
@77.2%
0.0%
0.216
@94.9%
0.199 0.142
GREG-mini 24
0.267
@98.1%
0.274
@99.4%
0.210
@94.3%
0.677
@99.2%
0.130
@99.8%
0.092 0.047
168
0.405
@91.6%
0.421
@86.4%
0.282
@77.2%
0.0%
0.188
@97.2%
0.156 0.074
JAP-mini 24
0.0%
1.457
@47.4%
0.0%
0.624
@100.0%
0.0%
0.275 0.234
168
0.0%
0.0%
0.0%
0.0%
0.0%
0.416 0.304
CAISO-mini 24
0.0%
0.421
@69.6%
0.0%
0.874
@48.9%
0.0%
0.138 0.170
168
0.0%
0.0%
0.0%
0.0%
0.0%
0.303 0.364
NYTS-mini 24
2.537
@95.7%
2.562
@97.1%
3.169
@96.3%
2.455
@99.9%
1.013
@100.0%
0.838 0.387
168
3.184
@81.8%
3.061
@82.4%
2.911
@90.0%
0.0%
1.138 @98.2% 1.643 0.700
BEAR-mini 12
0.0%
1.577
@63.6%
0.0%
1.209
@88.3%
0.0%
0.093 0.066
144
0.0%
0.0%
0.0%
0.0%
0.0%
0.540 0.211
The test results of multimodal forecasting models and LLMs (for multimodal forecasting) on Fidel-TS are shown in Table 5. The results show that less capable models like the Qwen series and Chattime struggle to benefit from textual information; in some cases, their multimodal performance is even worse than their unimodal, series-only counterpart (in Table 4). Although Deepseek-R1, the most capable in reasoning, performs well in some cases, it still falls short of smaller, specialized multimodal forecasting models in most cases.
This result stands in contrast to previously reported successes of LLMs in multimodal forecasting (Wang et al., 2024a; Williams et al., 2024). On one hand, prior benchmarks were often susceptible to causal and description leakage, allowing LLMs to effectively "cheat" by extracting ground truth from text rather than genuinely learning cross-modal relationships. On our leak-free benchmark, this shortcut is eliminated, thereby exposing their weakness in this core task. On the other hand, data scarcity in previous benchmarks led to simplistic, unrealistic tasks with sparse text and short prediction horizons. While LLMs performed well in these easy cases, our benchmark’s richer textual data and more practical, longer forecasting windows reveal a performance degradation, highlighting their struggles with complex, long-context scenarios.
Furthermore, the failure of LLMs extends beyond mere accuracy to critical issues of reliability. As documented in Table 4, LLMs already exhibit failures in unimodal forecasting due to output formatting errors or exceeding context length limits. Table 5 shows that the addition of text exacerbates this problem, leading to a higher rate of failure. These results demonstrate that the combination of more channels, longer prediction horizons and richer textual input creates a perfect storm of context explosion that LLMs cannot reliably handle.
While previous analysis highlights scenarios where multimodal inputs can underperform, it is equally crucial to examine their success cases. Table 5 shows a compelling pattern in which the multimodal forecaster FIATS consistently delivers the best performance across most datasets, outperforming unimodal and multimodal models substantially. This indicates that the textual information provided in our benchmark, when leveraged by an appropriate architecture, confers a tangible advantage.
Table 6: Performance comparison of FIATS with different textual inputs on BEAR-mini test set. All metrics are MSE.
Pred. Len.
+Ctrl. +Wea.
24 0.066 0.123
168 0.211 0.546
This observation motivates a deeper investigation to pinpoint the source of this performance gain. The BEAR dataset serves as a natural laboratory for this inquiry, as it uniquely provides two distinct types of textual data for the same time series: internal control events and external weather reports. This configuration allows us to move beyond simply confirming that text is helpful and instead ask "what kind of text drives performance". To this end, we conducted a targeted analysis comparing the performance of FIATS when supplied with each of these two textual inputs, as detailed in Table 6.
Our targeted analysis on the BEAR dataset provides a validation of our proposed Causal Soundness principle. Consistent with the results in Table 5, supplying the FIATS model with causally potent text from internal control events yields state-of-the-art. In contrast, weather data not only provides no benefit but also degrades performance compared to unimodal models, which is expected in an indoor, climate-controlled system where internal signals dominate as the primary causal drivers. This finding becomes more compelling when contrasted with the substantial performance gains from the same weather data on a causally-dependent system like NYTS (Table 5). This juxtaposition of results confirms our core principle: the efficacy of multimodal forecasting is not inherent to the auxiliary data but is dictated by the causal potency of that information relative to the target system’s specific dynamics. More results and analysis are summarized in Appendix E.2
5 Conclusion and Future Works
Conclusion
In this work, we addressed a critical flaw in the evaluation of time series forecasting models by formalizing the principles of high-fidelity benchmarking and introducing Fidel-TS, a new large-scale benchmark built upon them. Our comprehensive experiments validate this new paradigm, leading to two core insights: (1) our high-fidelity benchmark reveals a strikingly different model performance landscape from prior work, exposing significant biases in conventional benchmarks that stemmed from unrealistic evaluation setups and data leakage. (2) we demonstrate that causal relevance is a critical determinant of success in multimodal forecasting, where the value of auxiliary data is dictated by its causal connection to the target system. By providing a robust and causally-sound evaluation standard, Fidel-TS offers a clearer path for the community to develop the next generation of truly effective forecasting models.
Limitations and Future Works
To uphold the strict leak-free principles of our initial design, we deliberately focused on weather and control events as the primary textual modalities. We acknowledge that while these are verifiably causal, they may not be the most potent drivers in all scenarios. This presents a clear path for future research: expanding Fidel-TS to include richer, more complex textual sources such as economic reports or news archives. The primary research challenge will be to develop novel methodologies that can filter and align these noisier data streams while upholding the principles of high fidelity and causal soundness. The extensible architecture of our benchmark and framework is designed to facilitate such future explorations.
Ethic Statement

### v3

@@ 4.2 Experimental Setup
(1) Unimodal Forecasting Models, including MLP-based Dlinear (Zeng et al., 2023), FITS (Xu et al., 2024b), transformer-based PatchTST (Nie et al., 2023), iTransformer (Liu et al., 2024b), LLM-reprogramming-based GPT4TS (Zhou et al., 2023), and advanced foundation models Chronos (Ansari et al., 2024), TimeMoE (Shi et al., 2025) and Sundial (Liu et al., 2025), which pre-trained on vast numerical time series corpora.
(2) Multimodal Forecasting Models, including LLM-reprogramming-based GPT4MTS (Jia et al., 2024) and transformer-based FIATS (Xu et al., 2024a). They have different input requirements due to their distinct modeling approaches.
(3) LLMs, including general-purpose LLMs Qwen-2.5-14B-Instruct (Yang et al., 2024), Qwen-2.5-14B-Instruct-1M (Yang et al., 2025b), Qwen-3-14B (Yang et al., 2025a), Deepseek-R1 (DeepSeek-AI et al., 2025), and expert LLM Chattime (Wang et al., 2025) which fine-tuned based on Llama-2-7b (Touvron et al., 2023).
@@
Evaluation Metrics
We adopt the Mean Square Error (MSE) as the main evaluation metric:
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ (1)
We add a pass rate to evaluate the LLMs, defined identically to the pass@3 metric used for tasks like code generation or math problem-solving. This addresses cases where the LLM occasionally disregard prompt constraints, causing format errors (e.g., incorrect prediction length or variable numbers) and task failure.
Realistic Settings
To simulate diverse real-world time scales, our design aligns forecasting horizons with intuitive and operational periods. For instance, for datasets with an hourly frequency, we employ historical horizon of 360 (indicates half a month) and prediction lengths of {24, 168, 336, 720} which correspond to {1 day, 1 week, 2 weeks, and 1 month}. Our framework not only supports classic ratio-based splitting (e.g., 7:1:2) (Zhou et al., 2021; Wu et al., 2021), but also enables splitting by timestamps (e.g. Jan. 1st, 2021, Jan. 1st, 2022). This approach makes the evaluation setup more transparent and controllable, providing further safeguard against potential leakage issue. See Appendix D for more details of settings.
To guide our experiments, we formulate four Research Questions (RQs) and evaluate different types of models on various forecasting tasks.
4.3 RQ1: How do unimodal forecasting models perform on the new benchmark?
Table 3: Performance of unimodal forecasting models on the full test set of each dataset. All metrics are MSE. The best results are highlighted in bold. Domain-specific models are trained on the training set of each dataset. Foundation models are evaluated in zero-shot forecasting without fine-tuning on each training set.
Data -set Pred. Len. Domain-Specific Foundation Models
Patch. iTran. Dlin. FITS GPT 4TS Chro -nos Time -MoE Sun -dial
@@ 4.3 RQ1: How do unimodal forecasting models perform on the new benchmark?
To explore the ranking of the capabilities of the selected unimodal forecasting model on our benchmark, we evaluated all of them and summarized the results in Table 3. The main findings are: (1) Although most of the domain-specific models have previously claimed state-of-the-art status on classic unimodal benchmarks, the performance of them is highly dataset-dependent on our benchmark, with a different model achieving the best results on each dataset. (2) While large-scale foundation models claim to "fit all time series" and can outperform domain-specific models in zero-shot forecasting, their performance is merely comparable to specialized models on short forecast horizons and degrades significantly over longer ones. More results regarding the original sampling rate without downsampling are provided in Appendix E.1. All these suggest that the improvements observed on prior benchmarks are likely biased. On the new benchmark, no model consistently maintains a leading position, and the superior capabilities claimed by foundation models fail to materialize.
4.4 RQ2: How do models perform under the new generalization test?
@@
Table 4: Prior and new generalization test of domain-specific unimodal forecasting models. All metrics are MSE. We report the difference in MSE between generalization tests and in-domain evaluations in Table 3; a smaller MSE gap indicates a closer match between the underlying data distributions. The best results are highlighted in bold.
Dataset Pred. Len. Patch. iTrans. Dlinear FITS GPT4TS
NYTS-h-obs $\rightarrow$ NYTS-h-hid 24 0.611 (+0.010) 0.606 (+0.025) 0.650 (+0.012) 0.655 (+0.011) 0.603 (+0.021)
168 1.123 (+0.097) 1.142 (+0.101) 1.187 (+0.089) 1.228 (+0.103) 1.162 (+0.097)
336 1.310 (+0.134) 1.329 (+0.134) 1.374 (+0.112) 1.440 (+0.134) 1.325 (+0.131)
720 1.424 (+0.149) 1.445 (+0.150) 1.501 (+0.123) 1.602 (+0.145) 1.513 (+0.184)
CCP-h $\rightarrow$ NYTS-h-hid 24 0.835 (+0.234) 0.866 (+0.285) 0.758 (+0.120) 0.763 (+0.119) 0.867 (+0.285)
168 1.308 (+0.282) 1.670 (+0.629) 1.242 (+0.144) 1.281 (+0.156) 1.334 (+0.269)
336 1.468 (+0.292) 1.928 (+0.733) 1.438 (+0.176) 1.722 (+0.416) 1.681 (+0.487)
720 1.549 (+0.274) 2.130 (+0.835) 1.568 (+0.190) 1.890 (+0.433) 5.044 (+3.715)
Prior benchmarks often employ transfer strategies that force models to transfer across datasets from entirely different domains (e.g., ETT $\rightarrow$ Weather), or across subsets within the same domain but with mismatched sampling frequencies (e.g., ETTh1 $\rightarrow$ ETTm1) (Jin et al., 2024). However, our distinction between subjects and channels enables a more realistic transfer setup. For a selected dataset, we train models exclusively on the -obs subsets and evaluate them directly on the held-out -hid subsets, simulating a more practical scenario and avoiding cross-dataset transfer.
To find the reasonable method of generalization test, we conducted both the prior and new test strategies on NYTS dataset. As shown in Table 4, the cross-dataset setting (CCP $\rightarrow$ NYTS) results in higher MSE values, with GPT4TS exhibiting an error several times larger than other baselines at length 720. Additionally, the model rankings in this setting differ from the in-domain benchmarks. In contrast, our proposed strategy (NYTS-obs $\rightarrow$ NYTS-hid) yields lower losses and maintains relative rankings that align with in-domain results in Table 3. This indicates that our method of splitting by subject and channel provides a more reasonable evaluation of generalization capability by avoiding the large distribution shifts found in cross-dataset testing. More generalization test results can be found in Appendix E.2.
4.5 RQ3: Does multimodal data actually improve predictive performance?
To verify whether multimodal forecasting models can improve performance as they claim, we conducted a comparison and summarized the results in Table 5. The results show that performance gains are not observed uniformly across all models, directly indicating that prior benchmarks suffer from bias. Furthermore, to investigate the underlying causes of the performance improvements brought by multimodal information, we leveraged our benchmark to conduct a more in-depth analysis of both models and data.
Table 5: Performance of domain-specific multimodal forecasting models on full test set of each dataset. All the models are trained on the training set of each dataset. All metrics are MSE. We report the difference between the multimodal models and the best unimodal model (as shown in Table 3), red represents better performance (MSE decrease) and blue represents worse performance (MSE increase).
Model / Dataset GPT4MTS FIATS
24 168 336 720 24 168 336 720
CCP-h 0.162 0.212 0.226 0.236 0.109 0.159 0.168 0.183
GREG-h 0.093 0.139 0.151 0.169 0.044 0.065 0.082 0.094
JAP-h 0.358 0.543 0.605 0.656 0.333 0.451 0.500 0.502
CAISO-h 0.183 0.370 0.443 0.543 0.202 0.429 0.509 0.600
NYTS-h-obs 0.617 1.099 1.290 1.335 0.407 0.644 0.685 0.710
GPT4MTS FIATS
12 144 288 576 12 144 288 576
@@ 4.5 RQ3: Does multimodal data actually improve predictive performance?
(+Ctrl.)
0.177 0.591 0.874 1.096 0.094 0.222 0.298 0.323
BEAR-obs
(+Weather)
0.117 0.538 0.713 0.910 0.129 0.562 0.998 1.309
Model Comparison
GPT4MTS falls behind the best unimodal baseline as it restricts text usage to the historical horizon, treating it merely as a covariate that offers limited foresight. Conversely, FIATS achieves state-of-the-art performance by leveraging text within the prediction horizon, effectively using it as explicit state guidance to direct the forecasting trajectory. This difference demonstrates that the effectiveness of multimodal forecasting hinges on the model itself.
Ablation Study
The BEAR dataset provides an ideal experimental testbed for dissecting modality effectiveness, as it offers two distinct types of textual features: deterministic control events and stochastic weather reports. Results show that for the FIATS whose capabilities have already been proven by other datasets, incorporation of control events leads to a significant reduction in error, whereas the inclusion of weather data results in negligible improvements or even performance degradation. This discrepancy arises because control events are causally linked interventions that directly dictate system state changes, while weather variables exhibit a much weaker, indirect correlation with immediate system dynamics. In conclusion, the performance gains of multimodal forecasting depend on the causal relevance between textual information and the underlying time-series variables of the system.
4.6 RQ4: Are LLMs effective for forecasting tasks?
The rapid improvement in the general capabilities of LLMs has led to their evaluation on specialized time series forecasting tasks. Currently, both general and fine-tuned LLMs typically perform inference via direct prompting; for forecasting tasks, a standard approach involves constructing specific prompts based on forecasting scenario (Williams et al., 2024) (Wang et al., 2024a). To this end, we designed prompts for unimodal and multimodal contexts. Detailed prompt templates are provided in the Appendix F.1.
Table 6: Performance of LLMs on -mini test set. All metrics are MSE. The best results are highlighted in bold and the second-best are underlined. Considering the limitations of LLM context, we choose the univariate forecasting datasets and shorter prediction length (24, 168). "–" indicates that the filled-in prompt exceeds the context limit of the LLM.
Model / Setting CCP-h-mini GREG-h-mini NYTS-h-mini
24 168 24 168 24 168
Qwen2.5- 14B-Inst. 128K Uni. 0.311 0.865 0.229 0.497 1.884 3.599
Mul. 0.315 0.570 0.267 0.405 2.537 3.184
1M Uni. 0.289 0.374 0.213 0.279 1.893 2.378
Mul. 0.303 0.555 0.274 0.421 2.562 3.061
Qwen3 -14B 128K Uni. 0.235 0.298 0.153 0.263 2.236 2.546
Mul. 0.222 0.336 0.210 0.282 3.169 2.911
Deep.-R1 128K Uni. 0.182 0.267 0.113 0.213 1.071 2.154
Mul. 0.137 0.216 0.130 0.188 1.013 1.138
Chattime Uni. 0.263 0.457 0.223 0.486 0.998 2.312
Mul. 0.313 0.677 2.455
Model Comparison
The results in Table 10 yields two primary insights. First, among all evaluated LLMs, the state-of-the-art (SOTA) DeepSeek-R1 consistently achieves superior performance. The more advanced Qwen 3 also outperforms the previous generation Qwen 2.5 in most cases. This indicates that the performance ranking in forecasting tasks closely mirrors that of general-purpose capability, suggesting that robust general reasoning abilities translate effectively to time-series inference. Second, while the fine-tuned Chattime improves upon the Qwen Series and even outperforms DeepSeek-R1 in some cases, it remains unstable, particularly in multimodal scenarios. This leads to the conclusion that existing fine-tuning methods improves specific abilities but offers no stable forecasting advantage. Additional results for multivariate datasets are summarized in Appendix E.3.
Modality Comparison
For the Qwen Series and Chattime, the inclusion of multimodal information does not consistently translate into performance improvements; in many cases, the increased complexity and context length appear to burden the models, resulting in degraded predictive ability. Conversely, DeepSeek-R1 effectively leverages the additional semantic information, yielding superior results in multimodal scenarios compared to unimodal scenarios in the majority of test cases. This suggests that the utility of multimodal context is highly dependent on the model’s intrinsic reasoning capacity.
Case Study
To further investigate the source of the forecasting and reasoning capabilities of SOTA LLMs, we conducted a case study on representative examples (detailed in Appendix F.2). The study leads to the conclusion that when dealing with a unimodal scenario, DeepSeek-R1 focuses on statistical extrapolation, extracting historical trends and patterns to predict. When facing multimodal semantic context, DeepSeek-R1 use its logic-driven reasoning ability to analyze the relationship between historical weather and the time series to infer future dynamics based on provided future weather data.
Capability Concerns
Although LLMs have shown progress in forecasting tasks with the improvement of their capability, a broader comparison reveals a critical limitation: compared to the evaluation results of dedicated forecasting models on -mini subsets (in Appendix E.4), nearly all LLMs lag behind specialized baselines, while only DeepSeek-R1 surpass some of them. This suggests that previous benchmarks for evaluating the forecasting ability of LLMs may have offered overly simplified tasks with short contexts, or suffered from description leakage and temporal leakage that reduced the difficulty of prediction. In contrast, under our context-rich and leakage-free forecasting setting, LLMs encounter substantially greater challenges.
Reliability Concerns
Beside the forecasting error, we further evaluated the robustness of the LLMs given the observed failure of LLMs. The comprehensive pass@3 results are summarized in Appendix E.5. It demonstrates that while LLMs maintain high success rates in scenarios characterized by shorter prediction horizons and fewer variables, failure rates escalate substantially when the models are subjected to longer prediction length, multivariate forecasting tasks, or inputs enriched with extensive textual context. The results highlight the persistent insufficiency of current LLMs in handling long-context, long-term realistic forecasting tasks.
5 Discussion
5.1 Efficiency analysis
We summarized the parameter statistics for all models in Figure 2. For domain-specific models, linear models (FITS, DLinear) remain the most efficient, followed by transformer-based models where multimodal approaches (FIATS) surpass unimodal ones (PatchTST, iTransformer) in parameter size. For the LLM-based methods (GPT4TS/GPT4MTS), while total counts are higher, the trainable parameters remain limited due to the frozen GPT-2 backbone; The multimodal GPT4MTS is slightly larger than the unimodal GPT4TS. The parameter scale increases further with cross-domain pretrained Time Series Foundation Models (TimeMoE, Sundial, Chronos) and peaks with LLMs, ranging from SFT-based (Chattime) to the general LLMs (Qwen series and Deepseek-R1).
Figure 2: Comparison of model parameters
5.2 Limitations and Future Works
Data Perspective
To enforce a strict leak-free principle, we primarily use weather reports as textual inputs. Although weather is generally causally relevant, our ablation study shows that it is not always the most informative driver for every system. Future work will extend the dataset with more causally grounded textual sources while maintaining real-world fidelity to avoid leakage. This also calls for robust data filtering pipelines to ensure that models receive the most relevant contextual information for forecasting.
Model Perspective
We select representative models under practical computational constraints to reveal common challenges in long-term time-series forecasting. Given the rapid pace of architectural development, it is impractical to include all emerging methods, such as specific LLM prompting approaches (e.g., LLMP (Requeima et al., 2024), LLMTime (Gruver et al., 2023)). Future work can broaden these baselines to cover a wider range of models and techniques, further enhancing the long-term value of our benchmark.
6 Conclusion
In this work, we addressed a critical flaw in the evaluation of time series forecasting models by formalizing the principles of high-fidelity benchmarking and introducing Fidel-TS, a new large-scale benchmark built upon them. Our comprehensive experiments validate Fidel-TS as a rigorous paradigm, revealing a performance landscape different from prior biased benchmarks. The thorough evaluation of different categories of models reveals new insights.
7 Impact Statement